A patient is classified as having gestational diabetes if their average glucose level is above 140 milligrams per deciliter (mg/dl) one hour after a sugary drink is ingested. Rebecca's doctor is concerned that she may suffer from gestational diabetes. There is variation both in the actual glucose level and in the blood test that measures the level. Rebecca's measured glucose level one hour after ingesting the sugary drink varies according to the Normal distribution with μ= 148 mg/dl and σ = 9 mg/dl, Using the Central Limit Theorem, determine the probability of Rebecca being diagnosed with gestational diabetes if her glucose level is measured:

A. Once?

B. 10 times?

C. 12 times?

Sketch the required probability on a normal curve for each of parts 1a, 1b, and 1c. This means that you need 3 sketches with parameters labelled and the requisite probability shaded.
Answer #1

a)

P(X > 140)

Z = (X - mean)/sd = (X - 148)/9

P(X > 140)

= 0.813

phpNhUxTN.png

b)

Xbar follow normal distribution with mean = mu = 148 and sd = sigma/sqrt(n) = 9/sqrt(n)

here n = 10

hence

sd = 9/sqrt(10)= 2.846049

P(Xbar > 140)

= 0.9975

139.462 142.308 145 154 148 150.846 153.692 156.538

c)

n = 12

sd = 9/sqrt(12) = 2.5980762

P(Xbar > 140)

= 0.999
Answer #2

When \(\mathrm{n}=1\) :

Note that sample mean \(\bar{X}\) follows a Normal distribution with

Mean \(=\mu=148\)

Standard deviation \(=\sigma / \sqrt{n}=9\)

Note that \(z=(\bar{x}-\mu) /(\sigma / \sqrt{n})=(x-148) / 9 \quad\) - - (Eq.1) follows standard Normal.

When \(\bar{x}=140,\) from \((\) Eq. 1\(), z=-0.89\)

\(P(\bar{X}>140)=P(Z>-0.89)=1-P(Z<-0.89)=1-0.1867=0.8133\)

When \(\mathrm{n}=10\) :

Note that sample mean \(\overline{\mathrm{X}}\) follows a Normal distribution with

Mean \(=\mu=148\)

Standard deviation \(=\sigma / \sqrt{n}=2.846\)

Note that \(z=(\overline{\mathrm{x}}-\mu) /(\sigma / \sqrt{n})=(x-148) / 2.846 \quad\) - - -(Eq.1) follows standard Normal.

When \(\bar{x}=140,\) from \((\) Eq.1 \(), z=-2.81\)

\(\mathrm{P}(\overline{\mathrm{X}}>140)=\mathrm{P}(\mathrm{Z}>-2.81)=1-\mathrm{P}(\mathrm{Z}<-2.81)=1-0.0025=0.9975\)

When \(n=12\) :

Note that sample mean \(\bar{X}\) follows a Normal distribution with

Mean \(=\mu=148\)

Standard deviation \(=\sigma / \sqrt{n}=2.5981\)

Note that \(z=(\overline{\mathrm{x}}-\mu) /(\sigma / \sqrt{n})=(x-148) / 2.5981 \quad \ldots\) (Eq.1) follows standard Normal.

When \(\bar{x}=140,\) from \((\) Eq. 1\(), z=-3.08\)

\(P(\bar{X}>140)=P(Z>-3.08)=1-P(Z<-3.08)=1-0.001=0.999\)


answered by: Jamepuff